OpenStudy (anonymous):

!What is the value of log625^5 YOULL GET A MEDAL a. -4 b. -1/4 c. 1/4 d. 4

OpenStudy (jhannybean):

\[\log(625^5)\]First pull out the 5 out front. This is a log rule dealing with powers. \[5\log(625)\]now tell me, is 625 a perfect square?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

it is

OpenStudy (jhannybean):

and what is it's square?

OpenStudy (anonymous):

25

OpenStudy (jhannybean):

Good. So you now have \[5\log(25^2)\]We now do the same thing, can you tell me what will happen next if i use the power rule with this log function?

OpenStudy (anonymous):

it will me 625 well yea or no

OpenStudy (jhannybean):

where will the power of 2 go?

OpenStudy (jhannybean):

hitn: follow the first step i presented you with.

OpenStudy (anonymous):

on top of the 625 wouldnt it or 5

OpenStudy (jhannybean):

What are you saying...

OpenStudy (jhannybean):

Are you following what I am doing?

OpenStudy (anonymous):

sorry im in a hurry cus its midnight at im sleepy but i need to finish this so im tired

OpenStudy (jhannybean):

\[5\log(25^2)\] we got 25^2 because, as you mentioned, 25^2 = 625, hence why I asked you if it is a perfect square.

OpenStudy (anonymous):

yea

OpenStudy (jhannybean):

So as the power rule states, the 2 will come out front and multiply with the 5.\[5 \cdot 2 \log(25)\] 5 x 2 = 10 So we have \[10 \log(25)\]

OpenStudy (jhannybean):

25 is the perfect square of 5^2 so we can write log(25) as log(5^2)

OpenStudy (jhannybean):

So finally we will have \[10 \cdot 2 \log(5) =20\log(5) \]